# Extracting the Luttinger parameter from a single wave function

### v1

@@ Abstract
Introduction — The theory of Tomonaga-Luttinger liquid (TLL) is a great triumph of strongly correlated physics [1]. Based on previous work of Tomonaga [2], Luttinger introduced a model to study interacting fermions in one dimension (1D) [3]. It clearly exemplifies the peculiarity of reduced dimensionality because infinitesimal interaction in a 1D system drives it away from the Fermi liquid. In subsequent works [4, 5, 6, 7], the bosonization framework was established to provide a unified treatment for many strongly correlated problems in 1D. Spin chains are usually studied using the Jordan-Wigner transformation that converts spins to fermions. Gapless phases of bosons or fermions can be understood using free boson fields that correspond to density fluctuations. Instabilities of the TLL lead to symmetry-breaking ordered phases. It is not easy to realize 1D systems, but experimental investigations have been carried out in solid state and cold atom platforms [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Another context in which TLL thrives is the edge of certain two-dimensional topological states [18, 19, 20, 21].
@@
$H=\frac{v}{8\pi}\int^{L}_{0}\mathrm{d}x\,\left[(\partial_{x}\hat{\varphi})^{2}+(\partial_{x}\hat{\theta})^{2}\right].$ (1)
In the conformal field theory (CFT) language, this is a system of compactified bosons. Here $L$ is the length of the system, $v$ is the characteristic velocity, $\hat{\varphi}$ is a compactified boson field with radius $R$ (i.e., $\hat{\varphi}\sim\hat{\varphi}+2\pi R$), and $\hat{\theta}$ is the conjugate boson field of $\hat{\varphi}$ with radius $R^{\prime}=2/R$. The Luttinger parameter $K$ can be expressed using the radius as $K=R^{2}/4$. Its value is $1$ for free fermions and deviates from $1$ when interaction is turned on. For a given TLL, many properties are determined by $K$ so it is routinely computed in numerical studies. If the Hamiltonian has a microscopic U(1) symmetry (e.g., particle number conservation), $K$ can be obtained by comparing the ground-state energy of multiple systems with different U(1) charges [22]. This is feasible because $K$ can be interpreted as the stiffness of a superfluid. The powerlaw decay of certain correlation functions can be used to deduce $K$ when their bosonized expressions are known [1]. It has been found that $K$ also appears in the U(1) charge fluctuations of a subsystem [23, 24], the U(1) symmetry-resolved entanglement spectrum [25, 26], or the entanglement entropy associated with particle partition [27, 28].
In this Letter, we propose that the Luttinger parameter in microscopic models can be extracted from the eigenstate crosscap overlap (ESCO). It has three favorable features: i) a single wave function instead of multiple states is used; ii) data fitting or finite-size scaling is not needed; iii) microscopic U(1) symmetry is not required. To the best of our knowledge, no other method enjoys all three advantages simultaneously. Our method is motivated by classical results and recent progress in CFT and integrable models. When a CFT is defined on the Klein bottle, important information can be inferred from the thermal entropy [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. The Klein bottle can be viewed as a special cylinder with crosscaps at its ends. A simple circle would turn into a crosscap if all pairs of antipodal points are identified. It is possible to define a variety of crosscap states in the framework of boundary CFT [42]. For a rational CFT, the crosscap states have universal overlaps (“crosscap coefficients”) with its ground state [43, 44, 45]. In some integrable models, the crosscap states have been identified as “integrable boundary states” [46, 47, 48, 49, 50]. The overlaps between all eigenstates of these integrable models and the crosscap states can be expressed as determinants, but their physical significance is not transparent. We aim to establish a general framework for studying crosscap states in CFT (without and with perturbations) such that physical properties can be deduced from their overlaps with eigenstates.
@@
General construction of crosscap states — While crosscap is simple to visualize in the continuum, crosscap states in lattice models have only been constructed in a somewhat intuitive manner. The main obstacle is how to make a connection between the degrees of freedom in the continuum and lattice. To this end, we find it particularly instructive to use the modified Villain model [51]
$H_{\mathrm{mV}}=\sum_{j=1}^{N}\left[(2\pi R\,\hat{p}_{j})^{2}+{R^{\prime}}^{2}(\hat{X}_{j+1}-\hat{X}_{j}+2\pi\hat{n}_{j,j+1})^{2}\right]\,,$ (2)
@@
as illustrated in Fig. 1(a). It provides a specific lattice realization of the compactified boson Hamiltonian in Eq. (1). There are $N$ sites in the chain (labeled by $j$), the lattice spacing is $a$, and periodic boundary condition is imposed. $\hat{X}_{j}$ ($\hat{E}_{j,j+1}$) is the site (link) variable at site $j$ [link $(j,j+1)$] and $\hat{p}_{j}$ ($\hat{n}_{j,j+1}$) is its canonical momentum. The standard commutation relations $[\hat{X}_{j},\hat{p}_{l}]=[\hat{E}_{j,j+1},\hat{n}_{l,l+1}]=i\delta_{jl}$ are satisfied. The local Hilbert space at site $j$ [link $(j,j+1)$], denoted as $\mathcal{H}_{j}$ ($\mathcal{H}_{j,j+1}$), is spanned by the eigenbasis of $\hat{X}_{j}$ ($\hat{E}_{j,j+1}$) and expressed as $|X_{j}\rangle$ with $X_{j}\in\mathbb{R}$ ($|E_{j,j+1}\rangle$ with $E_{j,j+1}\in[0,2\pi)$). The full Hilbert space is thus the tensor product of local ones on all sites and links, $\mathcal{H}=\prod^{N}_{j=1}\mathcal{H}_{j}\otimes\mathcal{H}_{j,j+1}$. However, local operators $\hat{G}_{j}=e^{i(\hat{E}_{j,j+1}-\hat{E}_{j-1,j}-2\pi\hat{p}_{j})}$ commute with both terms in the Hamiltonian (2), $[\hat{G}_{j},\hat{p}_{j^{\prime}}]=[\hat{G}_{j},\hat{X}_{j^{\prime}+1}-\hat{X}_{j^{\prime}}+2\pi\hat{n}_{j^{\prime},j^{\prime}+1}]=0\;\forall j,j^{\prime}$, and hence also commute with the Hamiltonian, $[H_{\mathrm{mV}},\hat{G}_{j}]=0\;\forall j$. Thus, we may impose the Gauss law constraint $\hat{G}_{j}=1$ such that the model becomes a lattice gauge theory. The physical Hilbert space is the subspace $\mathcal{H}_{\mathrm{phy}}$ of the full space $\mathcal{H}$ that satisfies the Gauss law. The projection operator onto the subspace $\mathcal{H}_{\mathrm{phy}}$ is denoted as $\mathcal{P}_{G}$.
To transform the model from lattice to continuum, we introduce new variables $\hat{\theta}_{j}=R^{\prime}(\hat{X}_{j}+2\pi\sum_{k=1}^{j}\hat{n}_{k-1,k})$ and $\hat{\varphi}_{j,j+1}=R\hat{E}_{j,j+1}$ [51]. It is obvious that $[e^{i\hat{\theta}_{j}/R^{\prime}},\hat{G}_{j^{\prime}}]=[e^{i\hat{\varphi}_{j,j+1}/R},\hat{G}_{j^{\prime}}]=0\;\forall j,j^{\prime}$, so $e^{i\hat{\theta}_{j}/R^{\prime}}$ and $e^{i\hat{\varphi}_{j,j+1}/R}$ act within $\mathcal{H}_{\mathrm{phy}}$. The compactification radii of $\hat{\theta}_{j}$ and $\hat{\varphi}_{j,j+1}$ are $R^{\prime}$ and $R$, respectively. In terms of these new variables, the Hamiltonian in Eq. (2) becomes
$H_{\mathrm{mV}}=\sum_{j=1}^{N}\left[(\hat{\varphi}_{j+1,j+2}-\hat{\varphi}_{j,j+1})^{2}+(\hat{\theta}_{j+1}-\hat{\theta}_{j})^{2}\right]\,.$ (3)
The continuum limit is taken by sending $N\rightarrow\infty$ and $a\rightarrow 0$ while $L=Na$ is fixed as a constant. After the mapping $\hat{\varphi}_{j+1,j+2}-\hat{\varphi}_{j,j+1}\rightarrow a\partial_{x}\hat{\varphi}(x)$ and $\hat{\theta}_{j+1}-\hat{\theta}_{j}\rightarrow a\partial_{x}\hat{\theta}(x)$, Eq. (3) reduces to the TLL Hamiltonian (1).
Motivated by the geometric structure of crosscap, one should try to construct crosscap states by putting antipodal points into certain maximally entangled states. In the modified Villain model, this line of thought leads to two crosscap states
$|\mathcal{C}_{\pm}\rangle=\mathcal{P}_{G}\prod_{j=1}^{N/2}|\mathcal{B}^{\mathrm{V}}_{\pm}\rangle_{j,j+N/2}\otimes|\mathcal{B}^{\mathrm{L}}_{\mp}\rangle_{j,j+N/2}\,,$ (4)
where
$|\mathcal{B}^{\mathrm{L}}_{\mp}\rangle_{j,j+N/2}$ $=\int_{0}^{2\pi}\mathrm{d}E\,|E\rangle_{j,j+1}|\mp E\rangle_{j+N/2,j+N/2+1}\,,$
|±$|\mathcal{B}^{\mathrm{V}}_{\pm}\rangle_{j,j+N/2}$ $=\int_{-\infty}^{\infty}\mathrm{d}X\,|X\rangle_{j}|\pm X\rangle_{j+N/2}$ (5)
are generalized Bell pairs of antipodal links and sites, respectively. It can be verified that $e^{i(\hat{\theta}_{j}\mp\hat{\theta}_{j+N/2})/R^{\prime}}|\mathcal{C}_{\pm}\rangle=|\mathcal{C}_{\pm}\rangle$ and $e^{i(\hat{\varphi}_{j,j+1}\pm\hat{\varphi}_{j+N/2,j+N/2+1})/R}|\mathcal{C}_{\pm}\rangle=|\mathcal{C}_{\pm}\rangle$ for all $j$ (note that $e^{i\hat{\theta}_{j}/R^{\prime}}$ and $e^{i\hat{\varphi}_{j,j+1}/R}$ commute with $\mathcal{P}_{G}$). In the continuum limit, they become the “sewing conditions”
@@
$e^{i[\hat{\varphi}(x)\pm\hat{\varphi}(x+L/2)]/R}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,,$ (6)
that are imposed on the crosscap states by boundary CFT [52, 53, 54]. It is exhilarating that these conditions can be derived rigorously using our formalism.
For subsequent calculations, the crosscap states may be rewritten using the mode expansion [55]:
$\hat{\theta}(x)$ $=\tilde{\varphi}_{0}+\frac{4\pi x}{L}\pi_{0}+\sum_{k\neq 0}\frac{i}{k}\left[e^{i\frac{2\pi k}{L}x}a_{k}-e^{-i\frac{2\pi k}{L}x}\bar{a}_{k}\right],$
$\hat{\varphi}(x)$ $=\varphi_{0}+\frac{4\pi x}{L}\tilde{\pi}_{0}+\sum_{k\neq 0}\frac{i}{k}\left[e^{i\frac{2\pi k}{L}x}a_{k}+e^{-i\frac{2\pi k}{L}x}\bar{a}_{k}\right]\,,$ (7)
@@
where the zero modes and oscillatory modes satisfy $[\varphi_{0},\pi_{0}]=[\tilde{\varphi}_{0},\tilde{\pi}_{0}]=i$ and $[a_{k},a_{l}]=[\bar{a}_{k},\bar{a}_{l}]=k\delta_{k+l,0}$, respectively. $\varphi_{0}$ ($\tilde{\varphi}_{0}$) is an angular variable with radius $R$ ($R^{\prime}$), and $\pi_{0}$ ($\tilde{\pi}_{0}$) is the associated canonical momentum. The eigenvectors of $\pi_{0}$ and $\tilde{\pi}_{0}$ that can be annihilated by all bosonic annihilation operators $a_{k>0}$ and $\bar{a}_{k>0}$ are denoted as $|n,m\rangle$ ($n,m\in\mathbb{Z}$); they are highest weight states (Virasoro primary states) of the CFT and satisfy $\pi_{0}|n,m\rangle=\tfrac{n}{R}|n,m\rangle$ and $\tilde{\pi}_{0}|n,m\rangle=\tfrac{m}{R^{\prime}}|n,m\rangle$. Using the mode expansion, the compactified boson Hamiltonian (1) becomes
$H=\frac{2\pi v}{L}\left[\pi_{0}^{2}+\tilde{\pi}_{0}^{2}+\sum_{k=1}^{\infty}(a_{-k}a_{k}+\bar{a}_{-k}\bar{a}_{k})-\frac{1}{12}\right],$ (8)
@@
whose energy eigenstates are just Virasoro primary states $|n,m\rangle$ and the descendant states obtained by applying the creation operators $a_{k<0}$ and $\bar{a}_{k<0}$ on top of $|n,m\rangle$. The crosscap states $|\mathcal{C}_{\pm}\rangle$ can be written as
@@
$|\mathcal{C}_{-}\rangle$ $=\kappa_{-}\exp\left[-\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{m\in\mathbb{Z}}|0,2m\rangle$ (9)
with undetermined overall factors using the mode expansion [56]. The values of $\kappa_{\pm}$ cannot be fixed by normalization. Employing the “cut-and-sew” procedure, we can compare the partition function $Z^{\mathcal{C}}(L,\beta)=\langle\mathcal{C}|e^{-\beta H}|\mathcal{C}\rangle$ on a cylinder with two crosscap boundaries (“tree channel”) and the partition function $Z^{\mathcal{K}}(2\beta,L/2)$ on a Klein bottle (“loop channel”). This analysis yields $\kappa_{+}=\sqrt{R^{\prime}}$ and $\kappa_{-}=\sqrt{R}$ [56].
@@
An appealing property of the crosscap states is that their overlaps with certain highest weight states $|n,m\rangle$ in the CFT are universal numbers depending only on the compactification radius (equivalently, Luttinger parameter):
@@
$H_{\mathrm{XXZ}}=\sum_{j=1}^{N}\left(S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}+\Delta S^{z}_{j}S^{z}_{j+1}\right)$ (11)
with $\Delta\in(-1,1]$ for which $K=\frac{\pi}{2(\pi-\cos^{-1}\Delta)}$ [1]. The second one is the Haldane-Shastry (HS) model

### v2

@@ Abstract
Introduction — The theory of Tomonaga-Luttinger liquid (TLL) is a great triumph of strongly correlated physics [1]. Based on previous work of Tomonaga [2], Luttinger introduced a model to study interacting fermions in one dimension [3]. It clearly exemplifies the peculiarity of reduced dimensionality because infinitesimal interaction in a one-dimensional system drives it away from the Fermi liquid. In subsequent works [4, 5, 6, 7, 8], the bosonization framework was established to provide a unified treatment for many strongly correlated problems in one dimension. Spin chains are usually studied using the Jordan-Wigner transformation that converts spins to fermions. Gapless phases of bosons or fermions can be understood using free boson fields that correspond to density fluctuations. Instabilities of the TLL lead to symmetry-breaking ordered phases. It is not easy to realize one-dimensional systems, but experimental investigations have been carried out in solid state and cold atom platforms [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Another context in which TLL thrives is the edge of certain two-dimensional topological states [19, 20, 21, 22].
@@
$H=\frac{v}{8\pi}\int^{L}_{0}\mathrm{d}x\,\left[(\partial_{x}\hat{\varphi})^{2}+(\partial_{x}\hat{\theta})^{2}\right],$ (1)
which is a system of compactified bosons in the language of conformal field theory (CFT). Here $L$ is the length of the system, $v$ is the characteristic velocity, $\hat{\varphi}$ is a compactified boson field with radius $R$ (i.e., $\hat{\varphi}\sim\hat{\varphi}+2\pi R$), and $\hat{\theta}$ is the conjugate boson field of $\hat{\varphi}$ with radius $R^{\prime}=2/R$. The Luttinger parameter $K$ can be expressed using the radius as $K=R^{2}/4$. Its value is $1$ for free fermions and deviates from $1$ when interaction is turned on. For a given TLL, many properties are determined by $K$ so it is routinely computed in numerical studies. If the Hamiltonian has a microscopic U(1) symmetry (e.g., particle number conservation), $K$ can be obtained by comparing the ground-state energy of multiple systems with different U(1) charges [23]. This is feasible because $K$ can be interpreted as the stiffness of a superfluid. The powerlaw decay of certain correlation functions can be used to deduce $K$ when their bosonized expressions are known [1]. It has been found that $K$ also appears in the U(1) charge fluctuations of a subsystem [24, 25], the U(1) symmetry-resolved entanglement spectrum [26, 27], different types of entanglement entropy [28, 29, 30, 31, 32], and finite entanglement scaling analysis [33, 34].
In this Letter, we propose that the Luttinger parameter in microscopic models can be extracted from the eigenstate crosscap overlap (ESCO). It has three favorable features: i) a single wave function instead of multiple states is used; ii) data fitting or finite-size scaling is not needed; iii) microscopic U(1) symmetry is not required. To the best of our knowledge, no other method enjoys all three advantages simultaneously. Our method is motivated by classical results and recent progresses in CFT and integrable models. When a CFT is defined on the Klein bottle, important information can be inferred from the thermal entropy [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. The Klein bottle can be viewed as a special cylinder with crosscaps at its ends. A simple circle with all pairs of antipodal points identified becomes a crosscap. It is possible to define a variety of crosscap states in the framework of boundary CFT [48]. For a rational CFT, the crosscap states have universal overlaps (“crosscap coefficients”) with its ground state [49, 50, 51]. In some integrable models, the crosscap states have been identified as “integrable boundary states” [52, 53, 54, 55, 56]. The overlaps between all eigenstates of these integrable models and the crosscap states can be expressed as determinants, but their physical significance is not transparent. We aim to establish a general framework for studying crosscap states in CFT (without and with perturbations) such that physical properties can be deduced from their overlaps with eigenstates.
@@
General construction of crosscap states — While the crosscap is easy to visualize, the problem of what kind of states qualify as conformal crosscap states in CFTs has not been completely settled [49, 50, 51]. We take the viewpoint that a conformal crosscap state should satisfy two sets of constraints. First, the field operators on antipodal spatial coordinates are identified with each other in some way. If a CFT has Lagrangian description, this can be made precise using the path integral formalism [57, 58, 59]. These requisites are physical manifestations of the geometric structure of crosscap and referred to as “sewing conditions”. Second, the partition functions on a cylinder with crosscap boundaries (tree channel) and on a Klein bottle (loop channel) should be equal. This “loop-tree channel correspondence” allows one to determine certain coefficients in crosscap states. As we turn to lattice models, the problem becomes even more complicated because the lattice counterpart of a field operator is generally difficult to find. It is thus unlikely that a conformal crosscap state can be converted rigorously to a lattice one. Based on the intuitive picture of identifying antipodal spatial points, recent works have studied lattice crosscap states in which antipodal sites form Bell states [52, 53, 54, 55, 56], but their relation to the conformal crosscap states has not been elucidated.
It is particularly instructive to study a specific lattice realization of the Hamiltonian in Eq. (1), namely the modified Villain model  [60] as illustrated in Fig. 1(a)
$H_{\mathrm{mV}}=\sum^{N}_{j=1}\left[(2\pi R\,\hat{p}_{j})^{2}+{R^{\prime}}^{2}(\hat{X}_{j+1}-\hat{X}_{j}+2\pi\hat{n}_{j,j+1})^{2}\right].$ (2)
@@
There are has $N$ sites in the chain (labeled by $j$), the lattice spacing is $a$, and periodic boundary condition is imposed. $\hat{X}_{j}$ ($\hat{E}_{j,j+1}$) is the site (link) variable at site $j$ [link $(j,j+1)$] and $\hat{p}_{j}$ ($\hat{n}_{j,j+1}$) is the associated canonical momentum, and they satisfy the standard commutation relations $[\hat{X}_{j},\hat{p}_{l}]=[\hat{E}_{j,j+1},\hat{n}_{l,l+1}]=i\delta_{jl}$. The eigenbasis of these variables are denoted as $|X_{j}\rangle$ ($X_{j}\in\mathbb{R}$) and $|E_{j,j+1}\rangle$ [$E_{j,j+1}\in[0,2\pi)$], which span the local Hilbert space at site $j$ and link $(j,j+1)$, respectively. The tensor product of these local ones is the full Hilbert space. However, there are additional local operators $\hat{G}_{j}=e^{i(\hat{E}_{j,j+1}-\hat{E}_{j-1,j}-2\pi\hat{p}_{j})}$ that commute with the Hamiltonian. We define the physical subspace $\mathcal{H}_{\mathrm{phy}}$ by imposing the Gauss law constraints $\hat{G}_{j}=1\;\forall j$. The projection operator onto the subspace $\mathcal{H}_{\mathrm{phy}}$ is denoted as $\mathcal{P}_{G}$. Using new variables $\hat{\theta}_{j}=R^{\prime}(\hat{X}_{j}+2\pi\sum_{k=1}^{j}\hat{n}_{k-1,k})$ and $\hat{\varphi}_{j,j+1}=R\hat{E}_{j,j+1}$ [60], Eq. (2) can be rewritten as
$H_{\mathrm{mV}}=\sum^{N}_{j=1}\left[(\hat{\varphi}_{j+1,j+2}-\hat{\varphi}_{j,j+1})^{2}+(\hat{\theta}_{j+1}-\hat{\theta}_{j})^{2}\right]\,.$ (3)
Its continuum limit is taken by sending $N\rightarrow\infty$ and $a\rightarrow 0$ with $L=Na$ kept unchanged. After the substitutions $\hat{\varphi}_{j+1,j+2}-\hat{\varphi}_{j,j+1}\rightarrow a\partial_{x}\hat{\varphi}(x)$ and $\hat{\theta}_{j+1}-\hat{\theta}_{j}\rightarrow a\partial_{x}\hat{\theta}(x)$, the TLL Hamiltonian emerges from Eq. (3). The operators $e^{i\hat{\theta}_{j}/R^{\prime}}$ and $e^{i\hat{\varphi}_{j,j+1}/R}$ act within $\mathcal{H}_{\mathrm{phy}}$ thanks to the commutation relations $[e^{i\hat{\theta}_{j}/R^{\prime}},\hat{G}_{j^{\prime}}]=[e^{i\hat{\varphi}_{j,j+1}/R},\hat{G}_{j^{\prime}}]=0\;\forall j,j^{\prime}$. The compactification radii of $\hat{\theta}_{j}$ and $\hat{\varphi}_{j,j+1}$ are $R^{\prime}$ and $R$, respectively.
Motivated by the intuitive picture mentioned above and previous works [49, 50, 51, 52, 53, 54, 55, 56], we construct the Ansätze
$|\mathcal{C}_{\pm}\rangle=\mathcal{P}_{G}\prod^{N/2}_{j=1}|\mathcal{B}^{\mathrm{V}}_{\pm}\rangle_{j,j+N/2}\otimes|\mathcal{B}^{\mathrm{L}}_{\mp}\rangle_{j,j+N/2}\,$ (4)
for crosscap states by putting antipodal links and sites in the generalized Bell pairs
$|\mathcal{B}^{\mathrm{L}}_{\mp}\rangle_{j,j+N/2}$ $=\int^{2\pi}_{0}\mathrm{d}E\,|E\rangle_{j,j+1}|\mp E\rangle_{j+N/2,j+N/2+1}\,,$
|±$|\mathcal{B}^{\mathrm{V}}_{\pm}\rangle_{j,j+N/2}$ $=\int^{\infty}_{-\infty}\mathrm{d}X\,|X\rangle_{j}|\pm X\rangle_{j+N/2}\,,$ (5)
respectively. It can be verified that the relations
$e^{i(\hat{\theta}_{j}\mp\hat{\theta}_{j+N/2})/R^{\prime}}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,,$
$e^{i(\hat{\varphi}_{j,j+1}\pm\hat{\varphi}_{j+N/2,j+N/2+1})/R}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,$ (6)
are fulfilled for all $j$, which in the continuum limit becomes
@@
$e^{i[\hat{\varphi}(x)\pm\hat{\varphi}(x+L/2)]/R}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,.$ (7)
These identities are sewing conditions in the sense that field operators on antipodal spatial coordinates are combined to form certain operators under which the crosscap states are invariant. It is amusing that they can be derived rigorously using our formalism.
For subsequent calculations, we introduce the mode expansions [61]
$\hat{\theta}(x)$ $=\tilde{\varphi}_{0}+\frac{4\pi x}{L}\pi_{0}+\sum_{k\neq 0}\frac{i}{k}\left(e^{i\frac{2\pi k}{L}x}a_{k}-e^{-i\frac{2\pi k}{L}x}\bar{a}_{k}\right),$
$\hat{\varphi}(x)$ $=\varphi_{0}+\frac{4\pi x}{L}\tilde{\pi}_{0}+\sum_{k\neq 0}\frac{i}{k}\left(e^{i\frac{2\pi k}{L}x}a_{k}+e^{-i\frac{2\pi k}{L}x}\bar{a}_{k}\right),$ (8)
@@
where the zero modes and oscillatory modes satisfy $[\varphi_{0},\pi_{0}]=[\tilde{\varphi}_{0},\tilde{\pi}_{0}]=i$ and $[a_{k},a_{l}]=[\bar{a}_{k},\bar{a}_{l}]=k\delta_{k+l,0}$, respectively. $\varphi_{0}$ ($\tilde{\varphi}_{0}$) is an angular variable with radius $R$ ($R^{\prime}$), and $\pi_{0}$ ($\tilde{\pi}_{0}$) is the associated canonical momentum. The eigenvectors of $\pi_{0}$ and $\tilde{\pi}_{0}$ that can be annihilated by all bosonic annihilation operators $a_{k>0}$ and $\bar{a}_{k>0}$ are denoted as $|n,m\rangle$ ($n,m\in\mathbb{Z}$); they are highest weight states (Virasoro primary states) of the CFT and satisfy $\pi_{0}|n,m\rangle=\tfrac{n}{R}|n,m\rangle$ and $\tilde{\pi}_{0}|n,m\rangle=\tfrac{m}{R^{\prime}}|n,m\rangle$. The TLL Hamiltonian in Eq. (1) becomes
$H=\frac{2\pi v}{L}\left[\pi^{2}_{0}+\tilde{\pi}^{2}_{0}+\sum^{\infty}_{k=1}(a_{-k}a_{k}+\bar{a}_{-k}\bar{a}_{k})-\frac{1}{12}\right],$ (9)
@@
whose energy eigenstates are just Virasoro primary states $|n,m\rangle$ and the descendant states obtained by applying the creation operators $a_{k<0}$ and $\bar{a}_{k<0}$ on top of $|n,m\rangle$. The crosscap states can be expressed as
@@
$|\mathcal{C}_{+}\rangle$ $=\kappa_{+}\exp\left[\sum^{\infty}_{k=1}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{n\in\mathbb{Z}}|2n,0\rangle\,,$
$|\mathcal{C}_{-}\rangle$ $=\kappa_{-}\exp\left[-\sum^{\infty}_{k=1}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{m\in\mathbb{Z}}|0,2m\rangle.$ (10)
The overall factors are not fixed by normalization but should be chosen as $\kappa_{+}=\sqrt{R^{\prime}}$ and $\kappa_{-}=\sqrt{R}$ to fulfill the loop-tree channel correspondence [62].
The conformal crosscap states in Eq. (10) are scale-invariant (i.e., they contain no scale) and cannot be normalized. Their lattice counterparts may be normalizable [Eq. (14) for spin-1/2 chains defined below] or not normalizable [Eq. (4) for the modified Villain model] , depending on the local Hilbert space dimension. In fact, for a given lattice model whose low-energy theory is the compactified boson CFT, the crosscap states together with the coefficients $\kappa_{\pm}$ in Eq. (10) should be understood as the continuum description of their lattice counterparts $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$. This means that they are fully determined by the overlaps of lattice crosscap states with low-lying eigenstates in the continuum limit: $\langle\mathcal{C}_{\pm}|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle=\lim_{N\to\infty}\langle\mathcal{C}_{\mathrm{latt}}^{\pm}|\psi_{n,m;\{n_{k}\},\{\bar{n}_{k}\}}\rangle$, where $|\psi_{n,m;\{n_{k}\},\{\bar{n}_{k}\}}\rangle$ is the lattice counterpart of the CFT eigenstate $|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle$.
@@
An appealing property of $|\mathcal{C}_{\pm}\rangle$ is that their overlaps with certain highest weight states $|n,m\rangle$ are universal numbers depending only on the compactification radius (equivalently, the Luttinger parameter):
@@
$H_{\mathrm{XXZ}}=\sum^{N}_{j=1}\left(S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}+\Delta S^{z}_{j}S^{z}_{j+1}\right)$ (12)
with $K=\frac{\pi}{2(\pi-\cos^{-1}\Delta)}$ for $\Delta\in(-1,1]$ [1]. The second one is the Haldane-Shastry (HS) model